Operation Summary for October 28, 2009 - October 30, 2009 |
Wednesday Oct. 28 The day shift began with Operations monitoring the Tevatron (TeV) store 7294 (26.11E30) and the Recycler stash (372.3E10), with the Antiproton Source stacking (17.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7294 at 8:15 AM. FESS personnel cleaned the TeV RF pond water strainer. Operations put the TeV into shot setup and established store 7295 at 10:59 AM with an average initial luminosity (AIL) of 259.11E30. TeV personnel conducted separator scans between 1:30 and 3:10 PM. During the evening shift, the MI vacuum system front end needed to be rebooted twice. Cryo system personnel at CHL reported that their nitrogen re-liquefaction plant developed an oil leak and had to be shut down for repairs. Switchyard took some beam. Thursday Oct. 29 The midnight shift began with Operations monitoring store 7295 (79.75E30) and the stash (320.4E10), with Pbar stacking (20.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7295 at 2:54 AM. Operations put the TeV into shot setup and established store 7296 at 4:40 AM with an AIL of 276.13E30. Operations reported that problems with LRF2 and MIRF2, and a P1 transfer line beam permit trip had caused short downtimes. During the day shift, the TeV quenched at 9:57 AM when the cryo system at sector A4 lost its control power; store 7296 was lost. CDF and D0 accessed their collision halls and mechanical support personnel accessed the TeV. Booster personnel conducted damper studies. The NuMI horn developed a power problem. During the evening shift, TeV personnel conducted studies. Several CAMAC front ends and the TCLK front end required a reboot. Operations put the TeV into shot setup at 6:25 PM. The TeV suffered a sector E4 quench at 7:33 PM. At 11:04 PM, Operations put the TeV through a dry squeeze followed by a wet squeeze. Friday Oct. 30 The midnight shift began with Operations monitoring the stash (158.0E10), with Pbar stacking (10.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV into shot setup and established store 7306 at 2 AM with an AIL of 125.86E30. Operations reported that the Linac RF buncher, LRF2, BRF16, and a power supply (I:HT701) had caused short downtimes. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow TeV, Pbar, and Recycler studies. |
|
More Information Comments and Suggestions |
| last modified 11/02/2009 email Fermilab |
| Security, Privacy, Legal | |